OpenStudy (jazzyfa30):

help please find the exact values of each expression. A) sin45 degrees+ cos 60 degrees b)sin 45 degrees *cos 45 degrees c)2 sin pi/3-3 tan pi/6

OpenStudy (mathmale):

Good morning! Have you a table of values of the sine, cosine and tangent of "easy" angles such as 45 and 60 degrees? pi/6 and p/3 radians? If not, how would you evaluate the sine of 45 degrees?

OpenStudy (jazzyfa30):

i dont know what your talking about.

OpenStudy (jazzyfa30):

@thomaster

OpenStudy (mathmale):

Sorry, jazzyfa, but being able to find the sine, cosine and tangent of 30, 45, 60 and 90 degrees is absolutely basic to "knowing" trigonometry. Do you know what the name "sine" signifies? If so, how is the sine of an angle defined?

OpenStudy (jazzyfa30):

can you please just show me how to work the problem this has to be in in 30 mins

OpenStudy (mathmale):

Again, you'll need to be able to find the values of sin 45 degrees and cos 60 degrees, just to do the first problem. What is the value of sin 45? of cos 60?

OpenStudy (mathmale):

If you have a trig textbook, these values will be in a table in that book. If you're taking an online course, I would hope that your instructor has posted a similar table of values of the trig functions.

OpenStudy (jazzyfa30):

http://www.mathwords.com/t/trig_values_of_special_angles.htm im not in trig im in pre cal

OpenStudy (mathmale):

Good table. Please use this table to find sin 45 degrees. Also find cos 60 degrees.

OpenStudy (jazzyfa30):

sin 45 is \[\sqrt{2}/2\] and cos 60 is 1/2

OpenStudy (mathmale):

You're doing fine! Then: sin45 degrees+ cos 60 degrees=\[\frac{ \sqrt{2} }{ 2 }+\frac{ 1 }{ 2}=?\]

OpenStudy (mathmale):

Actually, this answer is good enough. I'd suggest you move on to the next question (B), and come back to (A) for more simplification, if you have time.

OpenStudy (jazzyfa30):

i have to finish it out all the way to the end

OpenStudy (mathmale):

Fine. Please explain to me what you mean by that and how you would "finish it all the way out." Better yet, do it.

OpenStudy (jazzyfa30):

idk all i know is that it has to be in simplest form

OpenStudy (mathmale):

I would either leave our result as is:\[\frac{ \sqrt{2} }{ 2 }+\frac{ 1 }{ 2},\] or combine the two fractions into one:\[\frac{ \sqrt{2} }{ 2 }+\frac{ 1 }{ 2}=\frac{ \sqrt{2}+1 }{ 2 }\]

OpenStudy (mathmale):

Now please do B.

OpenStudy (jazzyfa30):

sin 45 is squareroot 2/2 and cos 45 is squareroot 2/2

OpenStudy (mathmale):

Agreed. And what are you supposed to do with those two results?

OpenStudy (jazzyfa30):

combine

OpenStudy (mathmale):

"combine" refers to addition or subtraction. Is that what we have in B?

OpenStudy (jazzyfa30):

ohhh multiply

OpenStudy (mathmale):

Please do that multiplication, OK?

OpenStudy (jazzyfa30):

.5

OpenStudy (mathmale):

OK. I'd prefer you show your work, but yes, the correct result is 0.5. Part C? do as much of the work there as you can by yourself.

OpenStudy (jazzyfa30):

idk where to start

OpenStudy (mathmale):

1. find sin (pi/3) from your table. 2. multiply your result by 2 and simplify that result.

OpenStudy (jazzyfa30):

its 60 degrees

OpenStudy (mathmale):

Yes, and sin (pi/3) has what value?

OpenStudy (jazzyfa30):

its squareroot 3/2

OpenStudy (mathmale):

1. find sin (pi/3) from your table. 2. multiply your result by 2 and simplify that result.

OpenStudy (jazzyfa30):

1.7320

OpenStudy (mathmale):

True, but leave your answer as Sqrt(3):\[2\frac{ \sqrt{3} }{ 2 }=?\]

OpenStudy (jazzyfa30):

my answer is 0

OpenStudy (mathmale):

Right. Nice work. Good luck to you!

OpenStudy (jazzyfa30):

thanks

OpenStudy (mathmale):

Yo're welcome!